OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)?

OpenStudy (anonymous):

A) (x − 2)2 + (y + 3)2 = 5 B) (x + 2)2 + (y − 3)2 = 25 C) (x − 2)2 + (y + 3)2 = 25 D) (x − 2)2 − (y + 3)2 = 5